OpenStudy (anonymous):

1/2(2x^3y^4z^5)(3x^2y^6) kindly help me multiplying this equation?

OpenStudy (anonymous):

is it supposed to have the multiplication sign or ^ sign?

OpenStudy (anonymous):

\[1/2\left( 2x^{3}y ^{4}z ^{5} \right)\left( 3x ^{2}y ^{6} \right)=1/2\left( 6x ^{5}y ^{10}z ^{5} \right)\]

OpenStudy (anonymous):

the () serves as the multiplication sign

OpenStudy (anonymous):

vood can u give me ur solutions?

OpenStudy (anonymous):

it will be like how Vood has written it according to me

OpenStudy (anonymous):

my ans. is 3x^5y^10z^5

OpenStudy (anonymous):

how is that?

OpenStudy (anonymous):

just combine like terms..

OpenStudy (anonymous):

but im not so sure if i get the right ans...

OpenStudy (anonymous):

You made a mistake in there. Although you carried the 1/2 through the equation, notice that z^5 has no coefficients, as does the y^10 term. That means you cannot reduce it at all with a fraction, and it will end up being multiplied by it. A slightly more cumbersome, but correct, way of expressing the answer is 1/2(6x^5*y^10*z^5) = 3x^5(y^10*z^5)/2. I prefer to use the first eqn, because the second can be misinterpreted. But the 1/2 goes under the section in parantheses ONLY.